Nine years ago Jack was eleven times older than paul. Find Jacks present age if their combined age now is 90
Answer by ewatrrr(24785)


Previously Posted
x Paul's age now, y Jack's age now
x + y = 90 and y - 9 = 11(x-9) 0r 11x - y = 90
....
11x - y = 90
x + y = 90
12x = 180
x = 15, Paul's age now. Jack is 75 y+=+90-15